A fan blades start from rest and move with a constant angular acceleration and reaches an angular speed of 14.0rad/s.find (a) the magnitude of the angular acceleration of the wheel

(B) the angle (in radians) through which it rotates in this time.


1
Expert's answer
2022-01-31T14:23:32-0500

(a) the magnitude of the angular acceleration of the wheel:


"\alpha=\Delta \omega/t=(14-0)/t=14/t\text{ rad/s}^2."


Substitute time to get the answer.

(B) the angle (in radians) through which it rotates in this time:


"\phi=\omega t/2=7t\text{ rad}."


Substitute time to get the answer.